# Collaborative and parametric insurance on the Ethereum blockchain

### v1

@@ Abstract
This paper introduces a blockchain-based insurance scheme that integrates parametric and collaborative elements. A pool of investors, referred to as surplus providers, locks funds in a smart contract, enabling blockchain users to underwrite parametric insurance contracts. These contracts automatically trigger compensation when predefined conditions are met. The collaborative aspect is embodied in the generation of tokens, which are distributed to both surplus providers and policyholders. These tokens represent each participant’s share of the surplus and grant voting rights for management decisions. The smart contract is developed in Solidity, a high-level programming language for the Ethereum blockchain, and deployed on the Sepolia testnet, with data processing and analysis conducted using Python. In addition, open-source code is provided and main research challenges are identified, so that further research can be carried out to overcome limitations of this first proof of concept.
@@ 1 Introduction
A blockchain is a decentralized ledger maintained by consensus within a peer-to-peer network. The nodes in this network agree on a shared history of data by executing an algorithm known as a consensus protocol. The Bitcoin blockchain, as described by Nakamoto [22], employs an algorithm called proof-of-work (PoW) to validate and record transactions between users. In contrast, the Ethereum blockchain, introduced by Buterin [5], uses a proof-of-stake (PoS) mechanism. Ethereum is designed to operate as a "world computer," enabling developers to deploy compiled code to power decentralized applications (dApps) that users can interact with directly.
@@
Through these decentralized applications, Ethereum extends the capabilities of blockchain beyond simple fund transfers, enabling more complex financial operations and fostering the emergence of a new financial ecosystem known as decentralized finance (DeFi), as described by Schär [24]. Traditional finance relies on trusted third parties and centralized databases, which involve intermediary costs, lack transparency, operate only during limited hours, and are susceptible to cyberattacks. In contrast, distributed ledgers are publicly accessible, replicated in multiple locations, and function continuously, 24/7. Furthermore, the code that governs decentralized applications is auditable, improving transparency. For further discussion, we refer the reader to Lipton and Treccani [19, Chapter 2].
Some criticisms of the financial system can also be extended to the insurance sector. Insurance policies are often complex and difficult for the general public to understand. Premium calculations are typically based on historical data that are privately held by insurance companies, which limits transparency. In addition, the claim settlement process is frequently lengthy and requires on-site assessments by one or more experts. In response to these challenges, the industry has witnessed the emergence of a new type of insurance known as parametric or index-based insurance. For discussion of its application to crop insurance, see Abdi et al. [1] and Conradt et al. [6]. Parametric insurance provides compensation based on predefined triggers—such as specific weather conditions or flight delays—rather than relying on traditional, often time-consuming, claims assessments. Policyholders may, however, be exposed to basis risk—the risk that the index-based compensation does not accurately reflect the actual loss experienced by the policyholder. This straightforward mechanism can be easily encoded in smart contracts, making parametric insurance a promising candidate for blockchain-based insurance solutions.
We present a smart contract on the Ethereum blockchain to manage a portfolio of parametric insurance policies. Our framework is inspired by the discussion in Cousaert et al. [7], from which we adopt the role distribution, while introducing an additional mechanism for the "tokenization" of risk. The smart contract owner writes the code and sets the initial values for parameters such as the premium loading and the risk level required for risk capital allocation. Investors participate by locking funds in the smart contract, in exchange for which they receive protocol tokens. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. The contribution of each investor is reflected in their token holdings.
@@
Future policyholders interact with the protocol to seek protection against unfortunate events. An insurance agreement can only be concluded if the funds provided by the investors are sufficient to cover the risk. The smart contract funds must remain above a specified threshold at all times, as prescribed by standard insurance regulations such as the European directive Solvency II. A portion of the premiums paid by customers is used to reward them with protocol tokens, which can be exchanged for ETH at any time. The value of the token, or equivalently its exchange rate against ETH, fluctuates over time depending on whether compensation is paid.
The token ownership of policyholders incentivizes honest behavior. Token issuance through risk underwriting decreases the token’s value, encouraging investors to remain committed until certain insurance policies are resolved. Conversely, the token value increases if no compensation is paid at settlement. Additionally, token ownership enables governance of the insurance protocol. The contract owner may update the smart contract parameters based on recommendations from a board composed of token holders, including both investors and policyholders. This governance model transforms the smart contract into a Decentralized Autonomous Organization (DAO).
The tokenization of risk on the blockchain can be viewed as a novel way to design insurance-linked securities, as outlined in Barrieu and Albertini [4]. By enabling policyholders to become investors, our smart contract creates a collaborative insurance scheme, akin to those described in Feng [12]. Investors, referred to as Surplus Providers (SPs), are analogous to liquidity providers in decentralized exchanges, which serve as blockchain-based trading venues for crypto assets; see, for instance, Mohan [21].
@@
We aim to design our smart contract to operate with minimal human interference, maximizing automation. Every interaction with the smart contract incurs a cost, as any transaction submitted to the blockchain requires a fee to be processed by the network of nodes. Therefore, optimizing the number of transactions is crucial for reducing operational costs.
Actuarial methods for calculating risk capital are often too complex to be directly implemented in a high-level language like Solidity, which, for instance, only supports integer-based arithmetic. Additionally, deploying a smart contract on the blockchain entails costs that grow with its complexity, emphasizing the need for simplicity in design. One of our key contributions is the simplification of these computations through justified approximations. In particular, we propose a straightforward recursive formula to determine the solvency capital required to adequately cover the risks of a parametric insurance portfolio. Another contribution of this paper is the demonstration of the feasibility of deploying a beta version of such a smart contract, where certain complexities are intentionally simplified. We also identify the main limitations of this simplified framework and provide the research and actuarial communities with an open-source beta version of the smart contract. This prototype serves as a foundation for developing generalizations and improved versions that address these limitations.
@@
A smart contract is a system defined by state variables that evolve based on actions performed by users. Section 2 introduces the roles of the participants, including the contract owner, surplus providers, and policyholders. It also establishes a probabilistic framework for the state variables using stochastic processes. Examples of state variables include the smart contract’s balance and the tokens held by each participant. Users interact with the smart contract by invoking functions, and the possible actions and their impacts on the system’s state are detailed in Section 3. Section 4 defines the risk capital levels that the surplus should not fall below. It also outlines a method to compute the solvency capital required dynamically as new insurance contracts are added to the portfolio. Section 5 demonstrates how the smart contract operates through an event-driven scenario, showcasing all its functions and culminating in the liquidation and reset of the smart contract. The main limitations of this beta version of the smart contract are discussed in Section 6, along with potential ideas for addressing them.
@@ 2 Parametric insurance portfolio management framework
A smart contract defines a system whose state evolves based on the actions of blockchain users. This section describes the roles of the users of our smart contract in Section 2.1, followed by the definition of stochastic processes used to monitor the state of the system in Section 2.2.
2.1 Role distribution
The key participants in our blockchain insurance scheme are the smart contract owner, the investors also referred to as surplus providers, and the policyholders. Both the underwriting and claim assessment processes are fully automated within the smart contract protocol. Consequently, unlike in Cousaert et al. [7], our framework does not include claim assessors or underwriters. It is also worth noting that the role of our investors closely resembles that of reinsurers in Cousaert et al. [7].
2.1.1 Smart contract owner
The smart contract owner is responsible for designing the insurance protocol by writing the Solidity code for the smart contract. After deploying the smart contract on the Ethereum blockchain, the owner initializes two key parameters: the premium loading and the risk level associated with the solvency capital. The premium loading is divided into two components. The first component covers the management costs of the smart contract and provides returns for the investors. The second component is redistributed to the policyholders in the form of protocol tokens.
2.1.2 Surplus providers
Surplus providers, or investors, contribute financially to the smart contract to build solvency capital. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. In exchange for making their funds available, investors receive tokens that represent their share of the surplus. Surplus providers can withdraw a portion of their funds by burning tokens in exchange for an equivalent amount of ETH, calculated based on the current exchange rate. This exchange rate evolves over time, influenced by whether claims are being compensated. Withdrawals are permitted as long as the funds locked in the smart contract do not fall below a threshold known as the Solvency Capital Requirement (SCR), which corresponds to a quantile of the future loss distribution.
2.1.3 Policyholders
Policyholders seek protection against unfortunate events. If such an event occurs, the smart contract pays a lump-sum compensation denominated in ETH. To obtain this protection, policyholders pay a premium to the smart contract, also denominated in ETH. A small portion of this premium is redistributed to them in the form of protocol tokens. An insurance agreement can only be concluded if the value of the Solvency Capital Requirement (SCR), after including the new risk, does not exceed the current funds locked in the smart contract.
@@ 2.1.3 Policyholders
2.1.4 Decentralized Autonomous Organization
Token ownership grants the right to participate in the protocol’s governance. Token holders collectively form a board that can issue recommendations to modify the protocol, such as adjusting parameters like the safety loading or the level of risk. The responsibility for implementing these changes lies with the smart contract manager, who may also be one of the investors.
2.2 Mathematical framework
In this section, we introduce several stochastic processes to monitor the state of our smart contract. We first focus on the stochastic processes relevant to the management of the parametric insurance portfolio in Section 2.2.1. Then, we examine the stochastic processes associated with the balance of the smart contract and protocol token ownership in Section 2.2.2.
2.2.1 Parametric insurance portfolio over time
A parametric insurance contract $(A_{T},l)$ pays a predefined compensation $l$ if a specific event $A_{T}$ occurs at time $T>0$. Denote by $S$ the time at which the contract is underwritten. Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space such that the event $A_{T}$ is $\mathcal{F}_{T}$-measurable. Define the status of the contract over time as:
$\xi_{t}=\begin{cases}0,&\text{if }t<T,\\ 1,&\text{if }t\geq T\text{ and }\bar{A}_{T},\\ 2,&\text{if }t\geq T\text{ and }A_{T},\\ \end{cases}$ (1)
where the state space of $(\xi_{t})_{t\geq 0}$ is $\mathcal{S}=\{0,1,2\}$. Each state corresponds to a specific condition of the contract:
@@ 2.2.1 Parametric insurance portfolio over time
The smart contract may be forced to file for bankruptcy if the surplus falls below the minimum capital requirement at any point between $S$ and $T$. If this occurs, the policy is canceled, and the premium is refunded to the policyholder. In such cases, a fourth possible state could be introduced:
$3=\text{"canceled"}.$" .
However, to simplify the presentation, we do not consider this possibility at this stage.
Let $\theta=\mathbb{P}(A_{T})$ denote the probability of the event $A_{T}$. The pure premium associated with a contract $(A_{T},l)$ is given by
$\mu=\theta\cdot l,$
where $\theta=\mathbb{P}(A_{T})$. Let $S<T$ the time at which the policy is underwritten. The commercial premium is then given by
$\pi=(1+\eta_{S})\cdot\mu,$
using the expectation premium principle, where $\eta_{S}>0$ is the loading factor at time $S$ to ensure profitability on average and cover management costs. This loading is split into two components, $\eta_{S}=\eta_{S}^{(1)}+\eta_{S}^{(2)}$, where $\eta_{S}^{(1)}$ covers the protocol’s operational costs and rewards the investors, while $\eta_{S}^{(2)}$ is redistributed to the policyholder in the form of protocol tokens according to an exchange rate defined in Section 2.2.2. The safety loading may evolve over time depending on the decisions of the board of token owners. The processes $\left(\eta^{(1)}_{t}\right)_{t\geq 0}$, $\left(\eta^{(2)}_{t}\right)_{t\geq 0}$ and $\eta_{t}=\eta_{t}^{(1)}+\eta_{t}^{(2)},\text{ }t\geq 0$ are $\mathcal{F}_{t}$-adapted stochastic processes.

### v2

@@ Abstract
This paper introduces a blockchain-based insurance scheme that integrates parametric and collaborative elements. A pool of investors, referred to as surplus providers, locks funds in a smart contract, enabling blockchain users to underwrite parametric insurance contracts. These contracts automatically trigger compensation when predefined conditions are met. The collaborative aspect is embodied in the generation of tokens, which are distributed to surplus providers. These tokens represent each participant’s share of the surplus and grant voting rights for management decisions. The smart contract is developed in Solidity, a high-level programming language for the Ethereum blockchain, and deployed on the Sepolia testnet, with data processing and analysis conducted using Python. In addition, open-source code is provided and main research challenges are identified, so that further research can be carried out to overcome limitations of this first proof of concept.
@@ 1 Introduction
A blockchain is a decentralized ledger maintained by consensus within a peer-to-peer network. The nodes in this network agree on a shared history of data by executing an algorithm known as a consensus protocol. The Bitcoin blockchain, as described by Nakamoto [24], simply record transactions between bitcoin users. In contrast, the Ethereum blockchain, introduced by Buterin [6], is also equipped with a compiler to execute pieces of code called smart contract. Ethereum is designed to operate as a "world computer," enabling developers to deploy decentralized applications (dApps) that users can interact with directly.
@@
Through these decentralized applications, Ethereum extends the capabilities of blockchain beyond simple fund transfers, enabling more complex financial operations and fostering the emergence of a new financial ecosystem known as decentralized finance (DeFi), as described by Schär [27]. Traditional finance relies on trusted third parties and centralized databases, which involve intermediary costs, lack transparency, operate only during limited hours, and are susceptible to cyberattacks. In contrast, distributed ledgers are publicly accessible, replicated in multiple locations, and function continuously, 24/7. Furthermore, the code that governs decentralized applications is auditable, improving transparency. For further discussion, we refer the reader to Lipton and Treccani [20, Chapter 2].
Some criticisms of the financial system can also be extended to the insurance sector. Insurance policies are often complex and difficult for the general public to understand. Premium calculations are typically based on historical data that are privately held by insurance companies, which limits transparency. In addition, the claim settlement process is lengthy and costly as it requires on-site assessments by one or more experts.
One response to these challenges is parametric insurance, also known as index-based insurance. Parametric insurance provides compensation based on predefined triggers, such as specific weather conditions, rather than relying on traditional, often time-consuming, claims assessments. The main advantages include fast payouts and reduced management costs. When the verification of the triggering condition is done using a reliable and independent data source, it increases policyholders’ confidence in the system.
Parametric insurance is not a new concept. It emerged in Asia during the 1990s, allowing farmers to buy protection against income loss due to poor harvests following adverse weather conditions. Since then, parametric insurance has been used to cover perils like natural catastrophes. Notable examples include the Caribbean Catastroph Risk Insurance Facility (CCRIF), launched in 2007, and the African Risk Capacity (ARC) founded in 2014. Fast payouts in the context of natural disaster relief are crucial, as they enable the rapid deployment of resources by government agencies and non-governmental organizations providing immediate assistance.
In recent years, parametric insurance schemes have emerged in mass-market products, such as flight delay insurance offered by China Southern Airlines, which provides payouts of up to 300 yuan (approximately US\$40) for delays of three hours or more.
The main downside of parametric insurance is basis risk, where the insured’s losses may not match the coverage amount, or the parameter may not trigger. Effective structuring and pricing require understanding the policyholder’s exact risks and choosing the right parameter. The availability of finer granularity data and advances in predictive modeling can mitigate basis risk. Hybrid insurance solutions that combine parametric and standard indemnity mechanisms can also be effective, see for instance Lopez and Nkameni [21]. An initial payment following the parametric trigger can provide immediate relief, and this amount can later be deducted from the compensation resulting from the traditional claims assessment process. The policyholder has the option to either maintain parametric insurance or purchase an additional traditional insurance contract to cover basis risk. It is conceivable that some insurers might be willing to supplement parametric insurance in certain cases, as insurers and reinsurers often find themselves in this position. For instance, they might partially hedge their traditional insurance risk using a parametric Insurance-Linked Securities (ILS) solution, such as catastrophe bonds or mortality bonds. However, there has been a shift from parametric to indemnity triggers in the ILS sector, which means some challenges would need to be addressed. Note that death benefits in traditional insurance often consist of a fixed amount that varies according to limited number of parameters, such as the number of children or marital status. This forms a prominent example of a parametric insurance contract that is widely accepted and commercialized.
For a historical overview of parametric insurance, we refer the reader to the press article by Clere [7] in InsurTech Digital magazine. Interesting discussions on the use of parametric insurance to address climate risk issues can be found on the websites of the World Economic Forum [28] and the National Association of Insurance Commissioners [25]. The application of parametric insurance to crop insurance has been discussed in the literature (see Abdi et al. [1] and Conradt et al. [8]). The simplicity of the insurance mechanism makes it suitable for encoding in a smart contract. A blockchain can store data immutably and transparently, making the information associated with the triggering event and compensation available to all interested parties. Connecting the blockchain to users’ wallets allows for fast and automated processing of indemnity payments.
We present a smart contract on the Ethereum blockchain to manage a portfolio of parametric insurance policies. We have developed an actuarial framework suited for weather-linked triggering events, specifically monitoring daily rainfall. This framework establishes ground rules for underwriting, rate-making, and managing risk through the determination of solvency capital. Our smart contract is designed to operate with minimal human interference, maximizing automation. Every interaction with the smart contract incurs a cost, as any transaction submitted to the blockchain requires a fee to be processed by the network of nodes. Therefore, optimizing the number of transactions is crucial for reducing operational costs. Actuarial methods for calculating risk capital are often too complex to be directly implemented in a high-level language like Solidity, which, for instance, only supports integer-based arithmetic. Additionally, deploying a smart contract on the blockchain entails costs that grow with its complexity, emphasizing the need for simplicity in design. One of our key contributions is the simplification of these computations through justified approximations. In particular, we propose a straightforward recursive formula to determine the solvency capital required to adequately cover the risks of a parametric insurance portfolio.
The smart contract structure is inspired by the discussion in Cousaert et al. [10], from which we adopt the role distribution and the "tokenization" mechanism of risk. The smart contract owner writes the code and sets the initial values for parameters such as the premium loading and the risk level required for risk capital allocation. Investors participate by locking funds in the smart contract, in exchange for which they receive protocol tokens. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. The contribution of each investor is reflected in their token holdings. The value of the token, or equivalently its exchange rate against ETH, fluctuates over time depending on whether compensation are being paid to policyholders. Conversely, the token value increases if no compensation is paid at settlement. Token ownership enables governance of the insurance protocol. The contract owner may update the smart contract parameters based on recommendations from a board composed of token holders. This governance model transforms the smart contract into a Decentralized Autonomous Organization (DAO). The tokenization of risk on the blockchain can be viewed as a novel way to design insurance-linked securities, as outlined in Barrieu and Albertini [4]. By enabling policyholders to become investors, our smart contract creates a collaborative insurance scheme, akin to those described in Feng [12]. Investors, referred to as Surplus Providers (SPs), are analogous to liquidity providers in decentralized exchanges, which serve as blockchain-based trading venues for crypto assets; see, for instance, Mohan [23]. Pooling resources is not uncommon in the context of parametric insurance, as both the CCRIF and ARC are multi-country agreements to combat climate risk.
@@
Future policyholders interact with the protocol to seek protection against adverse events. An insurance agreement can only be concluded if the funds provided by the investors are sufficient to cover the risk. The smart contract funds must remain above a specified threshold at all times, as prescribed by standard insurance regulations such as the European directive Solvency II.
@@
This paper aims at showing the feasibility of deploying a beta version of such a smart contract, where certain complexities are intentionally simplified. We also identify the main limitations of this simplified framework and provide the research and actuarial communities with an open-source beta version of the smart contract. This prototype serves as a foundation for developing generalizations and improved versions that address these limitations.
@@
The rest of the paper is organized as follows. Section 2 provides a brief overview on blockchain technology. Section 3 presents an actuarial framework for managing parametric insurance contract portfolios. Section 4 focuses on the design of the smart contract. We introduce the participants, including the contract owner, surplus providers, and policyholders, in Section 4.1. A smart contract is characterized by state variables, which we describe in Section 4.2. Examples of state variables include the smart contract’s balance and the tokens held by each participant. Users interact with the smart contract by invoking functions, and the possible actions and their impacts on the system’s state are detailed in Section 4.3. Section 5 demonstrates how the smart contract operates through an event-driven scenario, showcasing all its functions and culminating in the liquidation and reset of the smart contract. The main limitations of this beta version of the smart contract are discussed in Section 6, along with potential ideas for addressing them.
2 Blockchain technology and smart contracts
The concept of blockchain technology was first introduced with Bitcoin, a decentralized digital currency launched in 2008 with Nakamoto [24]. Bitcoin’s blockchain is a public ledger that records all transactions in a secure, immutable, and transparent manner. Its core innovation lies in achieving consensus among distributed nodes without the need for a central authority, using a Proof-of-Work (PoW) mechanism.
However, the Bitcoin blockchain was designed with a narrow scope: it enables peer-to-peer transfer of value but offers very limited programmability. As a result, more complex applications, such as decentralized financial instruments or autonomous organizations, are not feasible on the Bitcoin network. To overcome these limitations, the concept of programmable blockchains emerged, most notably with the introduction of Ethereum in 2015, see Buterin [6]. Ethereum extends the blockchain model by embedding a Turing-complete virtual machine, the Ethereum Virtual Machine (EVM), which enables the execution of code on-chain. This innovation allows developers to write smart contracts—autonomous programs that run exactly as programmed—opening the door to a wide range of decentralized applications (dApps) across finance, insurance, governance or digital ownership, some concrete projects are mentionned in Table 1.
Category Project Description URL
Finance (DeFi) Uniswap A decentralized exchange (DEX) that allows users to swap ERC-20 tokens using liquidity pools and an automated market maker model. https://uniswap.org
Insurance Etherisc Provides decentralized parametric insurance, e.g., crop insurance for farmers, using smart contracts and weather data oracles. https://etherisc.com
Governance Aragon A platform for building and managing decentralized autonomous organizations (DAOs) with voting and fund management capabilities. https://aragon.org
Digital Ownership Decentraland A virtual reality platform on Ethereum where users can buy, develop, and monetize virtual plot of land using NFTs (Non-Fungible Tokens). https://decentraland.org
Table 1: Examples of decentralized applications across sectors.
"ERC-20 tokens" refer to fungible tokens on the Ethereum blockchain that are interchangeable and used for various transactions and financial operations, while "NFTs" represent unique digital assets, such as collectibles or in-game items, that are not interchangeable and verify ownership on the blockchain. Token generation is a central feature of the Ethereum blockchain as it enables the creation of diverse digital assets that power decentralized applications and facilitate transactions between users.
Ethereum is currently the most widely adopted programmable blockchain, and as such was selected for the implementation of our parametric insurance framework. Ethereum benefits from the largest and most active developer community among blockchain platforms. This has led to a wealth of tooling, documentation, libraries (e.g., Hardhat, Truffle, web3.js), and community support, which greatly facilitate the development and deployment of decentralized applications. Many formal verification tools and auditing services are tailored for Ethereum-based contracts, increasing trust and reducing the likelihood of critical vulnerabilities. The Ethereum blockchains successfully transited from Proof-of-Work to Proof-of-Stake inducing a significant reduction of its energy consumption and improved network scalability, helping to ensure long-term sustainability and compliance with environmental standards. Lastly, Ethereum hosts the majority of DeFi protocols, NFT platforms, and DAO frameworks. Its large user base and vibrant application ecosystem make it the natural choice for projects that require integration with or exposure to these decentralized systems. We acknowledge the existence of other programmable blockchains that offer valuable features. Binance Smart Chain (BSC) offers compatibility with Ethereum’s EVM but is more centralized, with fewer validators, potentially limiting trust and decentralization. Solana provides high throughput and low latency but suffers from frequent network outages and a more complex programming environment. Avalanche supports custom blockchains and fast finality, but its developer ecosystem and tooling are less mature compared to Ethereum.
@@ 3 Parametric insurance portfolio management framework
3.1 Parametric insurance contract over time
A parametric insurance contract pays a predefined compensation $l$ if some observable quantity $(Q_{t})_{t\geq 0}$ reaches a threshold $\overline{Q}$ at some time $T$. The quantity $Q:=(Q_{t})_{t\geq 0}$ is a stochastic process defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. A compensation is paid if the $\mathcal{F}_{T}$-measurable event: $\{Q_{T}>\overline{Q}\}$ occurs. A parametric insurance contract is represented as a 5-tuple $(S,T,Q,\overline{Q},l)$, where $S$ denotes the time at which the contract is underwritten. We outline in Remark 3.1 and Remark 3.2 two extensions associated to the time of the triggering event and the amount compensated respectively.
Remark 3.1.
When the trigger is based on weather condition, the insurance coverage usually concerns a time interval. The triggering event definition must be tailored to the meteorological quantity we monitor. We consider daily rainfall as an example later in the article and so a typical event would be
@@ 3.1 Parametric insurance contract over time
It is more convenient to treat $Q$ as a discrete time stochastic process or a time serie in that case. If we consider windspeed then time $T$ should be replaced by a time interval $(T_{-},T_{+})$ and the triggering event then becomes
$\{Q_{t}>\overline{Q}\text{, for some }t\in(T_{-},T_{+})\},\text{ or }\{Q_{t}>\overline{Q}\text{, for all }t\in(T_{-},T_{+})\},$
and $Q$ is a continuous time stochastic process.
Remark 3.2.
The compensation can be a function $l:Q_{T}\mapsto l(Q_{T})$ of the quantity $Q_{T}$. Here, we only consider a pre-specified lump sum payment.
Define the status of the contract over time as:
$\xi_{t}=\begin{cases}0,&\text{if }t<T,\\ 1,&\text{if }t\geq T\text{ and }Q_{t}\leq\overline{Q},\\ 2,&\text{if }t\geq T\text{ and }Q_{t}>\overline{Q}.\\ \end{cases}$ (1)
The state space of $(\xi_{t})_{t\geq 0}$ is $\mathcal{S}=\{0,1,2\}$. Each state corresponds to a specific condition of the contract:
@@
The smart contract may be forced to file for bankruptcy if the surplus falls below the minimum capital requirement at any point between $S$ and $T$. If this occurs, the remaining active policies are cancelled, and the premiums are refunded to the policyholders. This possibility is described in Section 4.3.4.
The pure premium associated with a contract $(S,T,Q,\overline{Q},l)$ is given by $\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})\cdot l$, where $\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})$ is the probability of the event $\{Q_{T}>\overline{Q}\}$ estimated at time $S$. The filtration $(\mathcal{F}_{t})_{t\geq 0}$ gathers all the available information at time $t\leq T$ relevant to the occurrence of event $\{Q_{T}>\overline{Q}\}$ or equivalently the study of the quantity $(Q_{t})_{t\geq 0}$. For weather-linked events, the quantity $Q_{T}$ may be forecasted based on the available information at time $t<T$ using weather forecast models. Concretely, the filtration $(\mathcal{F}_{t})_{t\geq 0}$ contains both observations and predictions available at time $t\geq 0$ and this must be accounted for in our actuarial pricing framework. Constraints on $S$ can be imposed to mitigate the risk of adverse selection as outlined in Remark 3.3.
@@
From a practical standpoint, we can assume the existence of a time $S^{\ast}=T-\Delta<T$ before which the triggering event cannot be reliably predicted. Therefore, the estimation of the probability of $\{Q_{T}>\overline{Q}\}$ relies solely on historical, observed data. If policyholders have access to a more accurate predictive model than that of the insurers, it could lead to adverse selection. A straightforward approach to mitigate this risk is to stipulate that a policy $(S,T,Q,\overline{Q},l)$ cannot be underwritten if $S>S^{\ast}$. We can assume that:
$\theta=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{s}),\text{ for any }s\leq S^{\ast}.$
We consider as a use case a parametric insurance coverage against rain episodes. An illustration is provided in 1 with a concrete estimation of the probability $\theta$.
Example 1.
Consider the insurance policy $(S,T,Q,\overline{Q},l)$, where $Q=(Q_{t})_{t\geq 0}$ is the daily precipitation height measured at the meteo station of Strasbourg airport (Entzheim). We assume that $S\leq S^{*}$. We consider a parametric model for the daily precipitation height where $Q_{T}|\mathcal{F}_{S}$ is distributed as a compound Poisson-gamma random variable. We have
$Q_{T}|\mathcal{F}_{S}\overset{\mathcal{D}}{=}\sum_{k=1}^{N_{T}}U_{k,T},$
where $N_{T}$ is a Poisson random variable $\mathsf{Poisson}(\lambda_{T})$ and the $U_{k}$ are iid gamma random variables $\mathsf{Gamma}(\alpha_{T},\beta_{T})$ independent from $N_{T}$. This model was proposed by Dunn [11] with $N_{T}$ corresponding to the number of rain episodes during a day and the $U_{k}$’s characterize their intensity. We assume that the parameters of the model $\lambda_{T}$, $\alpha_{T}$ and $\beta_{T}$ simply depend on the month associated with the date $T$. We calibrate this model using daily data from the years $2000$ to $2023$ provided by Météo France and retrieved from the data.gouv.fr. The time series is shown in Figure 1.
Figure 1: Daily rainfall measurements recorded at the STRASBOURG-ENTZHEIM station, spanning from January 1, 2000, to December 31, 2023.
The model is fitted using a simple partial method of moments so that $\widehat{\lambda}=-\log(\widehat{p}_{0})$, where $\widehat{p}_{0}$ is the proportion of days without rain episodes over a month. We further have
$\widehat{\alpha}=\frac{\mu^{2}}{\widehat{\lambda}\sigma^{2}-\mu^{2}}\text{ and}\widehat{\beta}=\frac{\mu}{\widehat{\lambda}\widehat{\alpha}}.$
This simple method is described for instance in Goffard et al. [15]. We have a compound Poisson-Gamma model for the STRASBOURG-ENTZHEIM station for each month; the estimations of the parameters allow us to estimate the number of rain episodes and their intensity for each month as shown in Figure 2.
Figure 2: Number of rain episodes per month and intensity of the rain episodes depending on the month in Strasbourg
The climate in Strasbourg is "semi-continental". In terms of precipitation, it is characterized by low precipitation levels during the winter months (replaced by snow) and few but intense rain episodes during the summer months. The model then provides us with the probability $\theta$ associated with the threshold $\overline{Q}$ for each month, see Figure 3.
Figure 3: Estimation of the probability $\mathbb{P}(Q_{T}>\overline{Q})$ depending on the month and the threshold $\overline{Q}\in\{0,3,5,10\}$
The estimation of $\theta$ yields the pure premium of the contract as $\theta\cdot l$.
Remark 3.4.
The pricing of our parametric insurance contract relies on a piecewise constant function $T\mapsto\theta(T)$. From a practical perspective, dealing with dates in Solidity (the programming language of the Ethereum blockchain) is not straightforward; furthermore, having to handle 12 clusters of dates is tedious. Our solution consists of writing $T$ as an integer number between 1 and 365 and approximating the function $T\mapsto\theta(T)$ through a polynomial. In our smart contract, we only consider one threshold with $\overline{Q}=5$. The polynomial approximation of degree 4 is provided in Figure 4.
Figure 4: Estimation of the probability $\mathbb{P}(Q_{T}>5)$ depending on the day of the year.
Concretely, we use Python to find the coefficients of the polynomial approximation:
$\theta(T)\approx a_{0}+a_{1}\cdot T+a_{2}\cdot T^{2}+a_{3}\cdot T^{3}+a_{4}\cdot T^{4},$
before hardcoding the formula in Solidity. We note in passing that Solidity could allow us to encode a pricing formula that could result from the fitting of a logistic regression model, which is suitable for modeling the probability of occurrence of an event, provided that the model does not include too many covariates. The use of logistic regression for ratemaking purposes of parametric insurance contracts has been explored in Figueiredo et al. [13].
Suppose that $S\leq S^{\ast}$, where $S^{\ast}$ has been defined in Remark 3.3. The commercial premium is then given by
$\pi_{S}=(1+\eta_{S})\cdot\theta\cdot l,$
using the expectation premium principle, where $\eta_{S}>0$ is the loading factor at time $S$ to ensure profitability on average and cover management costs. The safety loading evolves over time according to managerial decisions based on the available information $\mathcal{F}_{t}$. It constitutes an $\mathcal{F}_{t}$-measurable stochastic process denoted by $(\eta_{t})_{t\geq 0}$.
3.2 Parametric insurance portfolio over time
A parametric insurance portfolio is a collection of parametric insurance contracts $\{(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i})\}_{i\geq 1}$. These contracts may be associated with their own observable quantities $Q_{i}:=\left(Q_{t}^{i}\right)_{t\geq 0}$ defined on a common probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. Let us take a concrete situation in 2 by following up on 1.
Example 2.
In addition to Strasbourg, we wish to offer coverage against rain episodes in the city of Marseille. We assume that Strasbourg and Marseille are sufficiently far apart (see the map on Figure 5) so that the precipitation levels of these cities are not correlated.
Figure 5: Location of Marseille and Strasbourg in France
Each contract $i$ is then associated with a location so that $Q^{i}\in\{Q^{\text{Marseille}},Q^{\text{Strasbourg}}\}$ and a date $T_{i}$ for all $i\geq 1$. We assume that $Q^{\text{Marseille}}$ and $Q^{\text{Strasbourg}}$ are mutually independent. We assume that $S_{i}<S_{i}^{\ast}$ for each contract so as to model the precipitation height via a compound Poisson-gamma distribution at each location. The models are fitted to the same data as in 1 and result in two different risk profiles as illustrated on Figure 6, where we plot the probability $\mathbb{P}(Q_{T}>5)$ as a function of the month of $T$ for each location.
Figure 6: Estimation of the probability $\mathbb{P}(Q_{T}>5)$ depending on the month and the location.